5) Determine gypsum board required for residential building exterior walls, interior walls and ceiling area (20 points) Assum Not sure if my answers are correct please help, thanks! Construction quantity Surveying.
Answer #1

Hi,

Solution as follows:

Step 1: Calculation of Numbers of Gypsum Boards for interior and exterior walls:

Given Data:

· Wall gypsum board size = ½” x 4’ x 8’

· Ceiling height = 8 ft

· Length of exterior walls = 148 LF (One side board)

· Length of interior walls = 148 LF (Both side boards)

· Waste factor = 5%

Since height of wall is equal to length of gypsum board, each board can be fixed at wall without cutting at short edge. Hence,

148 Required boards for exterior walls = = 37 boards

Add waste @ 5% of 37 = 1.85

Total for exterior walls = 37 + 1.85 = 38.85 boards

2 x 148 Required boards for interior walls = = -= 74 boards

Add waste @ 5% of 74 = 3.70

Total for interior walls = 74 + 3.70 = 77.70 boards

Total requirement of wall gypsum boards = 38.85 + 77.70 = 116.55 boards say 117 boards of ½” thickness

Step 2: Calculation of Numbers of Gypsum Boards for ceiling:

Given Data:

· Wall gypsum board size = 5/8” x 4’ x 8’

· Ceiling Area= 24 ft x 50 ft

· Waste factor = 5%

Ceiling area 24 x 50 Required boards for ceiling = 7 = 37.5 boards way Area of one board - 4x8

Add waste @ 5% of 37.5 = 1.88

Total for ceiling = 37.5 + 1.88 = 39.38 boards say 40 boards of 5/8” thickness

Hope the solution provided above is clear. In case of any clarifications, please leave comment in comments section.

Please give us a good rating, if you are satisfied with the response.

How would you rate our reply?

Best Regards

Chegg Expert Team